OpenStudy (anonymous):

A rectangular garden has a width that is 8 feet less than twice the length. Fine the dimensions if the perimeter is 20 feet.

OpenStudy (anonymous):

For such problems, you should assign variable names to unknown quantities. Here, width and length are. So, call them W and L. Now, use the information given to write equations in terms of W and L and solve. There are two pieces of information given, which will result in two equations in two variables. You should be able to solve them.

OpenStudy (anonymous):

Width is 8 feet less than twice the length. W = ?

OpenStudy (anonymous):

8<2X

OpenStudy (anonymous):

IS THAT RIGHT

OpenStudy (anonymous):

W is the width. It is 8 less than twice the length. Twice the length is 2L. Width is 8 less (means you need to subtract 8 from 2L). So, W = 2L - 8. That is the first equation.

OpenStudy (anonymous):

Second equation comes from the perimeter information given. Perimeter equals twice the length plus twice the width. So, 2W + 2L = 20. That is the second equation. Now, you should solve the two.

OpenStudy (anonymous):

TO BE HONEST I DO CANT SOLVE IT

OpenStudy (anonymous):

DO U SUBSTITUTE THE W WITH THE 8

OpenStudy (anonymous):

W = 2L - 8 (equation 1) 2W + 2L = 20 (equation 2) So, you substitute W from equation 1 into equation 2. 2(2L - 8) + 2L = 20 4L - 16 + 2L = 20 6L = 36 L = 6 Now, plugging this value of L in the first equation, you get: W = 2*6 - 8 => W = 4

OpenStudy (anonymous):

THANKS GT